# Self-interacting dark matter interpretation of Crater II

### v1

@@ III.2 The stellar distribution
Figure 2: The evolution of the stellar half-light radius $R_{1/2}$ and the line-of-sight velocity dispersion of stars within $R_{1/2}$ for the benchmarks: SIDM $\sigma/m=10~{}{\rm cm^{2}/g}$ (left panel), $30~{}{\rm cm^{2}/g}$ (middle panel), and $60~{}{\rm cm^{2}/g}$ (right panel), as well as CDM in each panel for comparison. The stellar distribution follows an Einasto profile and the initial scale radii are $r_{E}=0.40~{}{\rm kpc}$ (E1; solid), $0.73~{}{\rm kpc}$ (E2; dashed), and $1.37~{}{\rm kpc}$ (E3; dotted); the circle, square, and diamond symbols denote their corresponding snapshots for evaluation, respectively. The data point with error bars are from the measurements of Crater II Caldwell et al. (2017).
Figure 3: The line-of-sight velocity dispersion profile of stellar particles at the initial (gray) and final (colored) snapshots for the benchmarks: SIDM $\sigma/m=10~{}{\rm cm^{2}/g}$ (left panel), $30~{}{\rm cm^{2}/g}$ (middle panel), and $60~{}{\rm cm^{2}/g}$ (right panel), as well as CDM in each panel for comparison. The stellar distribution follows an Einasto profile and the initial scale radii are $r_{E}=0.40~{}{\rm kpc}$ (E1; solid), $0.73~{}{\rm kpc}$ (E2; dashed), and $1.37~{}{\rm kpc}$ (E3; dotted). The data points with error bars are the measurements from Caldwell et al. (2017).
@@
Figure 4: Mass (left panel) and density (right panel) profiles of the dark matter (solid) and stellar (dashed) components for the CDM (black), SIDM $\sigma/m=10~{}{\rm cm^{2}/g}$ (blue), $30~{}{\rm cm^{2}/g}$ (magenta), and $60~{}{\rm cm^{2}/g}$ (red) benchmarks at the final snapshot. The stellar density of the simulated galaxies is normalized to match the observed central stellar density of Crater II. The cyan and green bands denote the $68\%$ credibility intervals for dark matter and stars, respectively, based on a fit to the kinematics of Crater II Caldwell et al. (2017). The data points in the left panel denote the enclosed masses evaluated within $R_{1/2}$ and $\frac{4}{3}R_{1/2}$ Caldwell et al. (2017).
We have seen that after tidal stripping both our SIDM and CDM cases can reproduce the enclosed mass within $r=R_{1/2}$ to be consistent with the measurements. Now we examine the stellar distribution of the simulated galaxies. As discussed in Section II, we used an Einasto profile in Equation 3 to model the initial stellar distribution with three values of the scale radius $r_{E}=0.40~{}{\rm kpc}$, $0.73~{}{\rm kpc}$, and $1.37~{}{\rm kpc}$ at $t=0~{}{\rm Gyr}$ as in Borukhovetskaya et al. (2022). We have further checked that the Einasto profile provide an excellent fit to the stellar density profile at later times as well and determined $r_{E}$ for each selected snapshot through the fit. Using the relation $R_{1/2}\approx 2.03r_{E}$, we convert $r_{E}$ to $R_{1/2}$ and compare it to the measured value of Crater II $R_{1/2}\approx 1~{}{\rm kpc}$. Moreover, we compute the line-of-sight velocity dispersion of the stellar particles.
@@
Figure 2 shows the evolution of the half-light radius $R_{1/2}$ and the line-of-sight velocity dispersion $\sigma_{\rm los}$ of star particles within $R_{1/2}$ for the benchmarks SIDM10 (left panel), SIDM30 (middle panel), and SIDM60 (right panel), as well as the CDM case in each panel for comparison. The three initial conditions $R_{1/2}\approx 2.03r_{E}=0.81~{}{\rm kpc}$, $1.48~{}{\rm kpc}$, and $2.78~{}{\rm kpc}$ are denoted as E1 (solid; dot), E2 (dashed; square), and E3 (dotted; diamond), respectively; for each case the symbol denotes the snapshot that we conduct the evaluation. For all simulated galaxies, their size converges from the different initial values after tidal evolution. Ref. Borukhovetskaya et al. (2022) shows this convergence behavior for CDM; here we find that the SIDM cases follow a similar trend, but with important distinctions.
In CDM, the galaxy size significantly reduces due to tidal truncation and all three cases with different initial $R_{1/2}$ values converge to $R_{1/2}\sim 0.3\textup{--}0.4~{}{\rm kpc}$ at the final snapshot, too small to be consistent with the measured one $R_{1/2}\approx 1~{}{\rm kpc}$ Torrealba et al. (2016); Caldwell et al. (2017). However, in SIDM, the size reduces mildly for the case with the largest initial $R_{1/2}$, while it actually increases for the one with the smallest initial $R_{1/2}$. Remarkably, for all SIDM benchmarks, the size converges at the final snapshot: $R_{1/2}\approx 1.1~{}{\rm kpc}$ for SIDM10 and SIDM60, and $R_{1/2}\approx 1.3~{}{\rm kpc}$ for SIDM30, in good agreement with the observed size of of Crater II.
@@
The upper limit of the galaxy size is largely set by $r_{\rm max}$. For the region beyond $r_{\rm max}$, stellar particles are stripped away significantly. From Figure. 1 (right panel), we see that $r_{\rm max}\sim 0.4~{}{\rm kpc}$ for CDM, while $r_{\rm max}\sim 1.5~{}{\rm kpc}$ for SIDM, at the final snapshot. Thus the latter can retain a larger galaxy size. Additionally, in SIDM the stellar distribution expands as the halo core forms Vogelsberger et al. (2014); Carleton et al. (2019); Yang et al. (2020). At the final snapshot, our simulated SIDM halos have a core size of $1~{}{\rm kpc}$, which is comparable to $R_{1/2}$ of the stars.
From Figure 2, we also see that $\sigma_{\rm los}$ of the stars within $R_{1/2}$ decreases more prominent in SIDM than in CDM, although the latter has smaller pericenter. At the final snapshot, $\sigma_{\rm los}\approx 4~{}{\rm km/s}$ and $3~{}{\rm km/s}$ for CDM and SIDM, respectively, and they are comparable to the measured value of Crater II $\sigma_{\rm los}=2.7\pm 0.3~{}{\rm km/s}$ Caldwell et al. (2017). This is expected as we purposely choose orbital parameters for each case to reproduce the measured mass of Crater II within $R_{1/2}$, which is in turn estimated using the measured $\sigma_{\rm los}$. What is remarkable is that without further adjustments SIDM can simultaneously predict a large galaxy size consistent with Crater II. This is not the case for CDM.
@@
In Figure 3, we further show the line-of-sight velocity dispersion profile of stellar particles at the initial and final snapshots for the benchmarks SIDM10 (left panel), SIDM30 (middle panel), and SIDM60 (right panel), as well as the CDM case in each panel for comparison. Despite the difference among three $\sigma_{\rm los}$ profiles initially, they converge at their final snapshot. The velocity dispersion increases in the region for $R\gtrsim 1.5~{}{\rm kpc}$ as more stars are escaping at large radii. In the inner region $R\lesssim 1.5~{}{\rm kpc}$, the stellar dispersion profile is almost flat in SIDM as the halo has a shallow density core. In CDM, $\sigma_{\rm los}$ increases towards the center of the galaxy as expected from a central density cusp.
III.3 The mass and density profiles
@@
Figure 4 shows mass (left panel) and density (right panel) profiles of the dark matter (solid) and stellar (dashed) components for our CDM and SIDM benchmarks at the final snapshot. For comparison, we also display the favored range of Crater II from Caldwell et al. (2017) (cyan: halo; green: stars; $68\%$ credibility intervals), based on the fit using a generalized NFW profile for the halo, allowing a density core. We again see that the SIDM benchmarks agree with the observations better than the CDM case. Overall, the CDM halo is too dense in the inner region, while the SIDM halos have a shallow density core with a size of $\sim 1~{}{\rm kpc}$, resulting in a more extended stellar distribution.
IV Discussion and Conclusions
@@
As we have shown, the exceptionally large size of Crater II implies that its halo has a $1~{}{\rm kpc}$ density core. In CDM, a core could form due to baryonic feedback Read and Gilmore (2005); Mashchenko et al. (2008); Pontzen and Governato (2012); Di Cintio et al. (2014); Chan et al. (2015). However, for a progenitor halo with mass $3\times 10^{9}~{}M_{\odot}$, similar to Crater II, the FIRE2 simulation shows the core size is ${\cal O}(10)~{}{\rm pc}$ Lazar et al. (2020), too small to play a role. Thus we conclude that baryonic feedback is unlikely to help reconcile CDM with the observations of Crater II.
For SIDM, there is a degeneracy between the cross section and the orbital parameters in determining the properties of satellite galaxies. Our three SIDM benchmarks are almost identical in the galaxy size and the stellar velocity dispersion at the final snapshot. To break the degeneracy, a precise determination of Crater II’s orbital parameters is needed. In fact, our Orbit 3 is based on Gaia Early Data Release 3 McConnachie and Venn (2020). For this orbit, we found that $\sigma/m\sim 60~{}{\rm cm^{2}/g}$ at $V_{\rm max}\sim 25~{}{\rm km/s}$ is needed to reproduce the observations of Crater II, while the tension with CDM becomes much more severe; see Fig. 1 (left panel; green dashed).
@@
Our SIDM interpretation of Crater II can be extended to Antalia II, another satellite with a large size $R_{1/2}\sim 2.9~{}{\rm kpc}$ and $\sigma_{\rm los}\approx 5.7~{}{\rm km/s}$ Torrealba et al. (2019). We expect that the self-interactions can produce a core in Antlia II. Furthermore, there are over $60$ confirmed or candidate satellite galaxies of the Milky Way, see, e.g., McConnachie (2012); Drlica-Wagner et al. (2020). They exhibit a great diversity in both size and enclosed mass within it, with $R_{1/2}$ ranging from $\sim 10~{}{\rm pc}$ to a few ${\rm kpc}$ and $V(R_{1/2})\sim 4$ to more than $10~{}{\rm km/s}$. Thus it is intriguing to test if SIDM can explain the full range of the diversity in the $V(R_{1/2})\textup{--}R_{1/2}$ plane of the satellite galaxies.
@@
Cosmological N-body simulations show that strong dark matter self-interactions can diversify inner dark matter densities of satellite halos Zavala et al. (2019); Turner et al. (2021); Correa et al. (2022); Yang et al. (2023c); Nadler et al. (2023), due to core formation and collapse Balberg et al. (2002); Koda and Shapiro (2011). For $\sigma/m\gtrsim 30\textup{--}50~{}{\rm cm^{2}/g}$ in ultra-faint dwarf halos, the collapse could occur within the Hubble time, yielding a high central density. The collapse timescale is extremely sensitive to the concentration $\propto c^{-7/2}_{200}$ Essig et al. (2019); Kaplinghat et al. (2019); Nadler et al. (2023); Zeng et al. (2022, 2023). The concentration of the Crater II halo is close to the median and hence it is in the core-expansion phase. Nevertheless, the dense satellites of the Milky Way, such as Draco, could be in the collapse phase Nishikawa et al. (2020); Sameie et al. (2020); Kahlhoefer et al. (2019); Correa (2021). Interestingly, $\sigma/m\sim 60~{}{\rm cm^{2}/g}$ on the Crater II mass scale is aligned with SIDM models proposed to explain the extreme diversity of dark matter distributions in other galactic systems, see, e.g., Yang et al. (2023c); Nadler et al. (2023).
@@
In summary, we have performed controlled N-body simulations and shown that the unusual properties of Crater II can be explained in SIDM. Dark matter self-interactions lead to core formation and boost tidal mass loss for the satellite halo, resulting in a low velocity dispersion of stars even if the halo has a relatively large pericenter. At the same time, the stellar distribution expands as the halo core forms and the galaxy size increases accordingly. In the future, we will expand our study to other satellite galaxies of the Milky Way and explore SIDM models that can explain the full range of the diversity of the satellites.
Acknowledgements.
@@
This work of H.-B.Y. and D.Y. was supported by the John Templeton Foundation under grant ID#61884 and the U.S. Department of Energy under grant No. de-sc0008541. The work of H.A. was supported in part by the National Key R&D Program of China under Grants No. 2023YFA1607104, 2021YFC2203100, and 2017YFA0402204, the NSFC under Grant No. 11975134, and the Tsinghua University Dushi Program. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the funding agencies.
Appendix
@@
Figure 5 shows the $R_{1/2}$ and $\sigma_{\rm los}(<R_{1/2})$ values for the SIDM60 benchmark with three different final snapshots denoted as $t_{1}$, $t_{2}$, and $t_{3}$. In the main text, we showed the properties of the simulated galaxies at the snapshot when they are well relaxed, i.e., $t_{3}$ for SIDM60. The main results do not change if we evaluate the properties at the three different final snapshots.
@@
Figure 5: Similar to the middle panel of Figure 2, but with three sequential final snapshots for the SIDM60 benchmark denoted as $t_{1}$, $t_{2}$, and $t_{3}$. The insert panel shows their corresponding locations in the $V(R_{1/2})\textup{--}t$ plane.

### v2

@@ 3.2 The stellar distribution
Figure 2: Left panel: the evolution of the stellar half-light radius $R_{1/2}$ and the line-of-sight velocity dispersion of stars within $R_{1/2}$ for SIDM60 O3 (red), CDM O3 (blue), and CDM O2 (black). The circle, square, and diamond symbols denote their corresponding snapshots after every pericenter passage for evaluation. Right panel: The line-of-sight velocity dispersion profile of stellar particles at the initial (gray) and final (colored) snapshots for SIDM60 O3 (red), CDM O3 (blue), and CDM O2 (black). For both panels, the stellar distribution follows an Einasto profile with three initial $R_{1/2}$ values $0.8~{}{\rm kpc}$ (E1; solid), $1.5~{}{\rm kpc}$ (E2; dashed), and $2.8~{}{\rm kpc}$ (E3; dotted); the data point with error bars are from the measurements of Crater II (Caldwell et al., 2017). The orbit O3 is consistent with the measurements of Gaia EDR3 (McConnachie & Venn, 2020; Ji et al., 2021; Pace et al., 2022), while the orbit O2 is designed such that the CDM halo can lose sufficient mass.
@@
Figure 3: Mass (left panel) and density (right panel) profiles of the dark matter (solid) and stellar (dashed) components for SIDM60 O3 (red), CDM O3 (blue), and CDM O2 (black) at the final snapshot. The stellar density of the simulated galaxies is normalized to match the observed central stellar density of Crater II. The cyan and pink bands denote the $68\%$ credibility intervals for dark matter and stars, respectively, based on a fit to the kinematics of Crater II (Caldwell et al., 2017). The data points in the left panel denote the enclosed masses evaluated within $R_{1/2}$ and $\frac{4}{3}R_{1/2}$ (Caldwell et al., 2017).
@@
Figure 2 (left panel) shows the evolution of the half-light radius $R_{1/2}$ and the line-of-sight velocity dispersion $\sigma_{\rm los}$ of the stellar particles within $R_{1/2}$ for SIDM60 O3 (red), CDM O3 (blue), and CDM O2 (black). The initial conditions $R_{1/2}=2.03r_{E}\approx 0.8~{}{\rm kpc}$, $1.5~{}{\rm kpc}$, and $2.8~{}{\rm kpc}$ are denoted as E1 (solid; dot), E2 (dashed; square), and E3 (dotted; diamond), respectively; for each case the symbols denote the evaluation at snapshots after every pericenter passage.
For CDM O2, the size converges to $R_{1/2}\approx 0.3\textup{--}0.4~{}{\rm kpc}$ from $R_{1/2}\approx 0.8\textup{--}2.8~{}{\rm kpc}$ after tidal evolution; see also Borukhovetskaya et al. (2022). We have checked that our final $R_{1/2}$ values agree with the estimates based on the tidal track of CDM satellite halos in Errani et al. (2022). The same trend applies to CDM O3, but the convergence is slower as it has weaker tides. It is clear that both CDM O2 and O3 cannot explain the observations of Crater II. The former’s stellar size is a factor $3$ smaller than Crater II’s $R_{1/2}\approx 1~{}{\rm kpc}$, while the latter’s velocity dispersion is a factor of $3$ higher than the measured one $\sigma_{\rm los}(<R_{1/2})\approx 2.7~{}{\rm km/s}$. Remarkably, SIDM60 O3 can simultaneously explain both stellar size and velocity dispersion of Crater II.
@@
The upper limit of the galaxy size is largely set by $r_{\rm max}$. For the region beyond $r_{\rm max}$, stellar particles are stripped away significantly. From Figure 1 (right panel), we see that $r_{\rm max}\sim 0.4~{}{\rm kpc}$ for CDM O2, while $r_{\rm max}\sim 1.5~{}{\rm kpc}$ for SIDM60 O3 at the final snapshot. Thus the latter can retain a larger galaxy size. More specifically, in SIDM there are two competing effects in affecting the size of the satellite galaxy: tidal truncation and SIDM core formation. The former tends to reduce the size, while the latter causes it to expand (Vogelsberger et al., 2014; Carleton et al., 2019; Yang et al., 2020), see also Appendix C. If the initial distribution of stars is diffuse ($R_{1/2}\approx 2.8~{}{\rm kpc}$), tidal truncation dominates and the final size is reduced. If it is compact ($R_{1/2}\approx 0.8~{}{\rm kpc}$), the halo core expansion drives the increase of the stellar size. For $\sigma/m\approx 60~{}{\rm cm^{2}/g}$, SIDM produces a halo core size of $1~{}{\rm kpc}$, which is comparable to $R_{1/2}$ of Crater II. We have also checked that the tidal radius of the SIDM60 halo is $1~{}{\rm kpc}$ at the final snapshot.
@@
In Figure 2 (right panel), we show the line-of-sight velocity dispersion profile of stellar particles at the initial (gray) and final snapshot for SIDM60 O3 (red), CDM O3 (blue), and CDM O2 (black). For each of the three cases, despite the difference among three initial $\sigma_{\rm los}$ profiles corresponding to E1 (solid), E2 (dashed) and E3 (dotted), the final $\sigma_{\rm los}$ profiles largely converge. For SIDM60, the profile is consistent with the measurements (Caldwell et al., 2017). It is almost flat for $R\lesssim 1~{}{\rm kpc}$ because the halo has a shallow density core, while $\sigma_{\rm los}$ increases in the region for $R\gtrsim 2~{}{\rm kpc}$, where more stars are escaping at large radii. For CDM O2, $\sigma_{\rm los}$ increases towards the center as expected from a central density cusp. For CDM O3, $\sigma_{\rm los}$ is too high to be consistent with the measurements. We have further checked the $\sigma_{\rm los}$ profile using data from Ji et al. (2021) and found it overall agrees with that from Caldwell et al. (2017).
3.3 The mass and density profiles
@@
Figure 3 shows mass (left panel) and density (right panel) profiles of the dark matter (solid) and stellar (dashed) components for the three cases at the final snapshot. For comparison, we also display the favored ranges of Crater II from Caldwell et al. (2017) (cyan: halo; pink: stars; $68\%$ credibility intervals), based on a fit using a generalized NFW profile for the halo, allowing a density core. We see that SIDM60 reproduces both halo and stellar profiles inferred from Crater II. This is not the case for the CDM halos.
4 Discussion and Conclusions
@@
As we have shown, the exceptionally large size of Crater II implies that its halo has a $1~{}{\rm kpc}$ density core. In CDM, a core could form due to baryonic feedback (Read & Gilmore, 2005; Mashchenko et al., 2008; Pontzen & Governato, 2012; Di Cintio et al., 2014; Chan et al., 2015). However, for a progenitor halo with mass $3\times 10^{9}~{}M_{\odot}$, similar to Crater II, the FIRE2 simulation shows the core size is ${\cal O}(10)~{}{\rm pc}$ (Lazar et al., 2020), too small to play a role. More work is needed to test if baryonic feedback can help reconcile CDM with the observations of Crater II.
@@
Our SIDM interpretation of Crater II can be extended to Antlia II, another satellite with a large size $R_{1/2}\sim 2.9~{}{\rm kpc}$ and $\sigma_{\rm los}\approx 5.7~{}{\rm km/s}$ (Torrealba et al., 2019), challenging CDM as well (Errani et al., 2022)Ji et al. (2021) found that there is a velocity gradient in Antlia II and a tentative one in Crater II. This may provide another test on the SIDM interpretation. Furthermore, there are over $60$ confirmed or candidate satellite galaxies of the Milky Way, see, e.g., McConnachie (2012); Drlica-Wagner et al. (2020). They exhibit a great diversity in both size and enclosed mass, with $R_{1/2}$ ranging from $\sim 10~{}{\rm pc}$ to a few ${\rm kpc}$ and $V(R_{1/2})\sim 4$ to more than $10~{}{\rm km/s}$. Thus it is intriguing to test if SIDM can explain the full range of the diversity in the $V(R_{1/2})\textup{--}R_{1/2}$ plane of the satellite galaxies.
@@
Cosmological N-body simulations show that strong dark matter self-interactions can diversify inner dark matter densities of satellite halos (Zavala et al., 2019; Turner et al., 2021; Correa et al., 2022; Yang et al., 2023a; Nadler et al., 2023), due to core formation and collapse (Balberg et al., 2002; Koda & Shapiro, 2011). For $\sigma/m\gtrsim 30\textup{--}50~{}{\rm cm^{2}/g}$ in ultra-faint dwarf halos, the collapse could occur within the Hubble time, yielding a high central density. The collapse timescale is extremely sensitive to the concentration $\propto(\sigma/m)^{-1}c^{-7/2}_{200}$ (Essig et al., 2019; Kaplinghat et al., 2019; Nadler et al., 2023; Zeng et al., 2022, 2023). The concentration of the Crater II halo is close to the median and hence it is in the core-expansion phase. Nevertheless, the dense satellites of the Milky Way, such as Draco, could be in the collapse phase (Nishikawa et al., 2020; Sameie et al., 2020; Kahlhoefer et al., 2019; Correa, 2021). Interestingly, $\sigma/m\sim 60~{}{\rm cm^{2}/g}$ on the Crater II mass scale is aligned with SIDM models proposed to explain the extreme diversity of dark matter distributions in other galactic systems, see, e.g., Yang et al. (2023a); Nadler et al. (2023). For the velocity-dependent SIDM model in Yang et al. (2023a), the effective cross section on average is $\sim 60~{}{\rm cm^{2}/g}$ at $V_{\rm max}\sim 5\textup{--}25~{}{\rm km/s}$, the relevant range for the Crater II halo.
@@
In summary, we have performed controlled N-body simulations and shown that the unusual properties of Crater II can be explained in SIDM. Dark matter self-interactions lead to core formation and boost tidal mass loss for the satellite halo, resulting in a low velocity dispersion of stars even if the halo has a relatively large pericenter. At the same time, the stellar distribution expands as the halo core forms and the stellar distribution correlates with the core size. In the future, we will expand our study to other satellite galaxies of the Milky Way and explore SIDM models that can explain the full range of the diversity of the satellites, as well as galaxies in the field (Gentile et al., 2004; Kuzio de Naray et al., 2010; Oman et al., 2015; Ren et al., 2019; Piña Mancera et al., 2022; Kong et al., 2022; Montes et al., 2024; Mancera Piña et al., 2024).
acknowledgments
@@
We thank Ethan Nadler and Laura Sales for comments on the manuscript. This work of H.-B.Y. and D.Y. was supported by the John Templeton Foundation under grant ID#61884 and the U.S. Department of Energy under grant No. de-sc0008541. The work of H.A. was supported in part by the National Key R&D Program of China under Grants No. 2023YFA1607104, 2021YFC2203100, and 2017YFA0402204, the NSFC under Grant No. 11975134, and the Tsinghua University Dushi Program. The opinions expressed in this publication are those of the authors and do not necessarily reflect the views of the funding agencies.
Appendix A Choice of the final snapshot
@@
Figure 4 (left panel) shows the $R_{1/2}$ and $\sigma_{\rm los}(<R_{1/2})$ values for SIDM60 O3 with three different final snapshots denoted as $t_{1}$, $t_{2}$, and $t_{3}$. For our main results, we showed the properties of the simulated galaxies at the snapshot when they are well relaxed, i.e., $t_{3}$ for SIDM60. The results remain largely the same if we evaluate the properties at the three different final snapshots.
@@
Figure 4: Left panel: Similar to the left panel of Figure 2, but with three sequential final snapshots for SIDM60 O3 denoted as $t_{1}$, $t_{2}$, and $t_{3}$. The insert panel shows their corresponding locations in the $V(R_{1/2})\textup{--}t$ plane. Right panel: Similar to the left panel of Figure 2, but for two additional SIDM simulations SIDM10 O5 (orange) and SIDM30 O7 (purple).
Appendix B Degeneracy
Figure 4 (right panel) shows the evolution of $\sigma_{\rm los}(<R_{1/2})\textup{--}R_{1/2}$ for two additional SIDM simulations SIDM10 O5 (blue) and SIDM30 O7 (magenta), where the parameters of the orbit O5 (O7) are $\mu_{\alpha^{*}}=-0.091~{}(-0.0807)$, $\mu_{\delta}=-0.187~{}(-0.148)$, and $r_{\rm p}\approx 19.9~{}(27.9)~{}{\rm kpc}$. We see that SIDM10 O5 and SIDM30 O7 could reproduce the stellar size and velocity dispersion of Crater II as SIDM60 O3. Thus there is a degeneracy effect between cross section and orbit. Note that orbits O5 and O7 are not directly motivated by the Gaia EDR3 measurements and their associated halos would be tidally disrupted after $\sim 8~{}{\rm Gyr}$.
Appendix C N-body simulations with live stellar particles
Figure 5: Left panel: Surface density profiles for SIDM60 O3 with live stellar particles, including the initial condition (solid), the one constructed from live stellar particles at the final snapshot (dot) and its fit to a Plummer profile (dotted), as well as the one constructed from dark matter particles using the tagging method (square) and its fit to a Plummer profile (dashed). Right panel: Final mass profiles for dark matter (solid) and stars constructed from live stellar particles (dotted) and dark matter particles using the tagging method (dashed) for SIDM60 O3. The cyan and pink bands denote the $68\%$ credibility intervals for dark matter and stars, respectively (Caldwell et al., 2017).
We further confirm the main results and the tagging method itself by performing additional N-body simulations with live stellar particles for SIDM60 O3. We assume the initial stellar distribution follows a Plummer profile as
$\rho(r)=\frac{3M_{\star}}{4\pi a^{3}}\left(1+\frac{r^{2}}{a^{2}}\right)^{-\frac{5}{2}},$ (C1)
where $M_{\star}$ is the total stellar mass and $a$ is the scale radius. We take $M_{\star}=3.37\times 10^{6}~{}M_{\odot}$ and $a=0.71~{}{\rm kpc}$. For the Plummer profile, the 2D half-light radius is $R_{1/2}=a$, hence the initial value of $R_{1/2}$ is $0.71~{}{\rm kpc}$, which is comparable with the E1 initial condition for the Einasto profile, i.e., $R_{1/2}\approx 2.03r_{E}=0.81~{}{\rm kpc}$. We take the Plummer profile for the stars as it is relatively straightforward to implement in the code SpherIC (Garrison-Kimmel et al., 2013). For the dark matter halo, we take $\rho_{s}=1.02\times 10^{7}~{}M_{\odot}/{\rm kpc^{3}}$ and $r_{s}=2.30~{}{\rm kpc}$. The concentration of this halo is slightly lower than the one without stars to offset the baryonic influence of live stellar particles. This is a subtle and novel effect of SIDM and we will elaborate it further in a future publication.
Figure 5 (left panel) shows the initial (gray) and final (red) surface density profiles for SIDM60 O3 from N-body simulations with live stellar particles. We construct the stellar surface density profile at the final snapshot from stellar particles (dot), fit it with a Plummer profile (dotted), and determine the half-light radius at the final snapshot $R_{1/2}=a\approx 1.35~{}{\rm kpc}$. For comparison, we also use the tagging method to reconstruct the surface density (square), fit it with a Plummer profile (dashed), and find $R_{1/2}\approx 1.2~{}{\rm kpc}$. We see that $R_{1/2}$ at the final snapshot is increased almost by a factor of $2$ compared to the initial condition, due to SIDM core formation. Furthermore, $R_{1/2}$ inferred from the tagging method agrees with that from live stars within $11\%$, although the former underestimates the overall normalization of the surface density by a factor of $3$. We use the tagging method for the purpose of tracking the evolution of the galaxy size, and we can add a normalization factor to match the observed surface density. Furthermore, we have checked that for SIDM60 O3, the size evolution is almost the same whether using Einasto or Plummer profiles in the tagging method.
We also note that the surface density decreases in the central regions more significantly than in the outer regions. This is because SIDM core formation leads to a shallower density profile and weaker gravitational potential. Thus, the stellar particles more likely migrate from the center towards the outer regions and compensate for the loss there, compared to the CDM limit.
Figure 5 (right panel) shows the final mass profiles for dark matter (solid) and stars constructed from live stars (dotted) and dark matter particles using the tagging method (dashed). We see that the SIDM60 simulation with live stellar particles well reproduces the inferred halo and stellar mass profiles of Crater II (Caldwell et al., 2017).